Question1

  1. Describe and illustrate the expected relationship between risk and expected return based on security market line (SML)             (7mark)
  2. Hwang DBS Bhd has identified three potential stocks to buy in FBM KLCU but will only pick two to be added in the portfolio. The following information has been collected:  

    AirAsia

    Daya Material

    Muhibbah Engineering

    Expected return

    12%

    25%

    18%

    Standard deviation of return

    2.2

    5.9

    3.4

Covariances of returns between companies

AirAsia and Daya Material

8.2

AirAsia and Muhibbah Engineering

2.5

Daya Material and Muhibbah Engineering

17.6

(i)Estimate the correlation coefficients between each of the three possible portfolio combinations and explain the implications of these coefficients for portfolio risk.    (9mark)

(ii)Calculate the expected rate of return and standard deviation for each of the three possible 2-stock portfolios, and indicate which portfolio has the most efficient risk-return profile.                 (14mark)
Answer #1

a) The security market line is used in CAPM, a model that describes risk-return relationships for E(R) Undervalued Stocks Seci. Let A be AirAsia, B be Daya Material, C be Muhibbah Engineering CORR(RA,R3) = 7 - COV(RARB) (SD(RA) SD(RB)) 1 = 0.632 (2.2For second possible portfolio (B,C) E(Rp) =W3 X E(R3) + W X E(R) E(Rp) = 0.50 x 0.25 +0.50 x 0.18 E(Rp) = 0.215 Risk can be c